# Dense Video Object Captioning from Disjoint Supervision

### v2

@@ 0.B.3 Detailed captioning results
Table 9: Detailed captioning metrics of our zero-shot evaluation (Tab. References). We show the individual captioning metrics CIDEr, METEOR, and SPICE for each row on both datasets.
@@
Table 10: Detailed captioning metrics of our finetuning evaluation (Tab. References). We show the individual captioning metrics CIDEr, METEOR, and SPICE for each row on both datasets.
@@
STGRN [83] - - - 38.0 48.5 19.8
STVGBert [56] - - - 47.3 - 24.0
TubeDETR [72] 56.4 47.2 28.7 59.0 48.1 30.4
STCAT [31] - - - 61.7 50.8 33.1
@@
0.B.4 VidSTG spatio-temporal grounding evaluation
Table References of the main paper compared to prior methods on the spatial-video grounding task on VidSTG (where the input videos were assumed to already be temporally trimmed). In Tab. 11, we report results for spatial-, temporal- and spatio-temporal grounding by reporting the Spatial IoU (sIoU), Temporal IoU (tIoU) and Video IoU (vIoU) respectively.
@@ 0.B.4 VidSTG spatio-temporal grounding evaluation
Our model was designed for the Dense VOC task, and not grounding, and we were able to perform grounding by selecting the bounding boxes with the highest likelihood of generating the target sentence (Sec. References). As shown in Tab. 11, this approach works well for spatial-grounding, outperforming prior works in terms of the sIoU. However, as our model first generates object trajectories, without taking the input query into account, it struggles more at temporal localization. Nevertheless, it still achieves competitive results compared to prior works for both the tIoU and vIoU, although our model was not designed specifically for this task like the other methods in Tab. 11 which include explicit temporal localization modules within the network.
Appendix 0.C Broader impact and potential negative impact
Our work presents a new task and model for dense video object captioning. This task represents a general technology with a wide range of potential applications. Whilst we are unaware of all potential applications of such models, it is important to be cognizant that each application has its own merits and societal implications depending on the intentions of the individuals building and using the system. For example, we believe that the Dense VOC models can be used as part of systems to improve video search and retrieval, though they could also be used in video surveillance systems too. Additionally, we note that training datasets, especially for captioning [28, 84], can contain biases that may render models trained on them unsuitable for deployment.

### v3

@@ C.4 Detailed captioning results
Table 11: Detailed captioning metrics of our zero-shot evaluation (Tab. 3 left). We show the individual captioning metrics CIDEr, METEOR, and SPICE for each row on both datasets.
@@
Table 12: Detailed captioning metrics of our finetuning evaluation (Tab. 3 right). We show the individual captioning metrics CIDEr, METEOR, and SPICE for each row on both datasets.
@@
STGRN Zhang et al. (2020) - - - 38.0 48.5 19.8
STVGBert Su et al. (2021) - - - 47.3 - 24.0
TubeDETR Yang et al. (2022) 56.4 47.2 28.7 59.0 48.1 30.4
STCAT Jin et al. (2022) - - - 61.7 50.8 33.1
@@
C.5 VidSTG spatio-temporal grounding evaluation
Table 7 of the main paper compared to prior methods on the spatial-video grounding task on VidSTG (where the input videos were assumed to already be temporally trimmed). In Tab. 13, we report results for spatial-, temporal- and spatio-temporal grounding by reporting the Spatial IoU (sIoU), Temporal IoU (tIoU) and Video IoU (vIoU) respectively.
@@ C.5 VidSTG spatio-temporal grounding evaluation
Our model was designed for the Dense VOC task, and not grounding, and we were able to perform grounding by selecting the bounding boxes with the highest likelihood of generating the target sentence (Sec. 3.5). As shown in Tab. 13, this approach works well for spatial-grounding, outperforming prior works in terms of the sIoU. However, as our model first generates object trajectories, without taking the input query into account, it struggles more at temporal localization. Nevertheless, it still achieves competitive results compared to prior works for both the tIoU and vIoU, although our model was not designed specifically for this task like the other methods in Tab. 13 which include explicit temporal localization modules within the network.
Appendix D Limitations
Currently, our model produces a single caption for each trajectory, and in future work, we aim to caption potentially multiple action segments within a trajectory. Also, we repurposed existing grounding datasets for our task, as annotating a new captioning dataset can be subjective. We leave annotating a Dense VOC dataset with rigorous protocols and richer captioning as a future work.
Appendix E Broader impact and potential negative impact
Our work presents a new task and model for dense video object captioning. This task represents a general technology with a wide range of potential applications. Whilst we are unaware of all potential applications of such models, it is important to be cognizant that each application has its own merits and societal implications depending on the intentions of the individuals building and using the system. For example, we believe that the Dense VOC models can be used as part of systems to improve video search and retrieval, though they could also be used in video surveillance systems too. Additionally, we note that training datasets, especially for captioning Hendricks et al. (2018); Zhao et al. (2021), can contain biases that may render models trained on them unsuitable for deployment.